OpenStudy (anonymous):

Use the substitution method to solve x^2+2y^2=9 x-y+3=0 can someone show me the steps?

OpenStudy (anonymous):

I know the answer is x=1 and y=2 just let me get you the steps

OpenStudy (anonymous):

okay let's call the equations (1) and (2) respectively solve (2) for y, you can easily get: y=x+3 (3) now substitute the for y in equation (1): \[x^2+2(x+3)^2=9 \implies x^2+2(x^2+6x+9)=9\] \[x^2+2x^2+12x+18=9 \implies 3x^2+12x+9=0\] divide by 3 for simplification and factorize the expression \[x^2+4x+3=0 \implies (x+1)(x+3)=0\] \[x=-1 , x=-3\] substitute in equation (3) to get the correspondent values of y to each value of x